OpenStudy (anonymous):

Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8).

OpenStudy (anonymous):

... y=mx+b y being your y cordinate x being your x b being your y intercept and m being your slope... or change. when your giving two points like that u just have to find the change. so its pretty much Y2-Y1/X2/X1= the slope. so its 10-8/5-(-1) 2/6 which is 1/3 that means ur line changes for ever 1 rise u move to the right 3. y=1/3x+(look for your y intercept.) and that should be ur equation.

OpenStudy (anonymous):

There are several ways to go about this. The most basic is to first determine the slope using the points given. y2-y1/x2-x1 (8-10)/(5--1)= -2/6 which reduces to -1/3 This is the slope or m Next use the equation y=mx+b to determine b. Substitute values. I'm going to use the point (5,8) 8=-1/3(5)+b 8=-5/3+b 8+5/3=-5/3+5/3+b 29/3=b Rewrite in slope intercept form y=mx+b or y=-1/3x+29/3

OpenStudy (anonymous):

Studenttrainee, I think you dropped a negative in the slope

OpenStudy (anonymous):

10-8/-1-5 =2/-6 or -1/3

OpenStudy (anonymous):



OpenStudy (anonymous):

right